OpenStudy (anonymous):

Given f(x) = the quantity of x minus 5, divided by 3, solve for f-1(3)

OpenStudy (anonymous):

the equation would be f(x)=(x-5)/3

OpenStudy (anonymous):

thank you! can i ask another?

OpenStudy (anonymous):

do you know how to solve the equation? and yes

OpenStudy (anonymous):

would the answer be 21?

OpenStudy (anonymous):

im confused as to what f is so i dont know

OpenStudy (anonymous):

it says

OpenStudy (anonymous):

Given f(x) = the quantity of x minus 5, divided by 3, solve for f-1(3). -1 14 18 21

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

So we are given that \(f(x) = y = \dfrac{x-5}{3}\) To find \(f^{-1}(x)\), just swap x and y to each other then solve for y. So we have \(x = \dfrac{y-5}{3}\)

geerky42 (geerky42):

Can you solve for y? @dontsmileurugly

OpenStudy (anonymous):

no :S

geerky42 (geerky42):

You need to islate y. take all of these number to other side, where x is. Maybe start by multiplying both sides by 3?

geerky42 (geerky42):

isolate y*

OpenStudy (anonymous):

so 15 and 9

geerky42 (geerky42):

Multiply both sides by 3, what do you get?

OpenStudy (anonymous):

15

geerky42 (geerky42):

How?

OpenStudy (anonymous):

5x3

geerky42 (geerky42):

\(x=\dfrac{y-5}{3}\) \(x\color{red}{*3}=\dfrac{y-5}{3}\color{red}{*3}\) \(3x = \cdots?\)

OpenStudy (anonymous):

@dontsmileurugly when you multiply 3 it cancels out the other 3

OpenStudy (anonymous):

so its just 3?

OpenStudy (anonymous):

3x=y-5